\begin{thebibliography}{99}
\label{bibl}

%\bibitem[AP1]{AP1} {\sc A.B. Aleksandrov} and {\sc V.V. Peller}, {\em Functions  of perturbed operators},
% C.R. Acad. Sci. Paris, S\'er I {\bf347} (2009), 483--488.

\bibitem[AP1]{AP2}  {\sc A.B. Aleksandrov} and {\sc V.V. Peller},  {\em Operator H\"older--Zygmund functions}, Advances in Math.
{\bf224} (2010), 910-–966.

%\bibitem[AP3]{AP3}  {\sc A.B. Aleksandrov} and {\sc V.V. Peller},  {\em Functions of operators under perturbations of
%class $\bS_p$}, J. Funct. Anal. {\bf258} (2010), 3675--3724.
%
%\bibitem[AP4]{AP4}  {\sc A.B. Aleksandrov} and {\sc V.V. Peller},
%{\em Functions of perturbed unbounded self-adjoint operators.
%Operator Bernstein type inequalities}, Indiana Univ. Math. J. {\bf59:4}
%(2010), 1451 -- 1490.
%
%\bibitem[AP5]{AP5}  {\sc A.B. Aleksandrov} and {\sc V.V. Peller},
%{\em Functions of perturbed dissipative operators}, St.Petersburg Math. J. {\bf23:2} (2011), 9--51.
%
%\bibitem[AP6]{AP6}  {\sc A.B. Aleksandrov} and {\sc V.V. Peller}, {\it Estimates of operator moduli of continuity},
%J. Funct. Anal.

\bibitem[AP2]{AP}  {\sc A.B. Aleksandrov} and {\sc V.V. Peller}, {\sc Operator Lipschitz functions}, to appear.

%\bibitem[APPS1]{APPS} {\sc A.B. Aleksandrov, V.V. Peller, D. Potapov}, and
%{\sc F. Sukochev}, {\em Functionsof perturbednormal operators}, C.R. Acad. Sci. Paris, S\'er I {\bf348} (2010), 553--558.
%
%\bibitem[APPS2]{APPS2} {\sc A.B. Aleksandrov, V.V. Peller, D. Potapov}, and
%{\sc F. Sukochev}, {\em Functions of normal operators under perturbations},
%Advances in Math. {\bf226} (2011), 5216-–5251.

%\bibitem[Be]{Be} {\sc G. Bennett}, {\em Schur multipliers}, Duke Math. J. {\bf44} (1977), 603--639.

\bibitem[BS1]{BS1} {\sc M.S. Birman} and {\sc M.Z. Solomyak},
{\em Double Stieltjes operator integrals},
Problems of Math. Phys., Leningrad. Univ. {\bf1} (1966), 33--67 (Russian).
English transl.: Topics Math. Physics {\bf1} (1967), 25--54, Consultants Bureau Plenum
Publishing Corporation, New York.

\bibitem[BS2]{BS2} {\sc M.S. Birman} and {\sc M.Z. Solomyak},
 {\em Double Stieltjes operator integrals. II},
 Problems of Math. Phys., Leningrad. Univ. {\bf2} (1967), 26--60 (Russian).
English transl.: Topics Math. Physics {\bf2} (1968), 19--46, Consultants Bureau Plenum
Publishing Corporation, New York.

\bibitem[BS3]{BSsl} {\sc M.S. Birman} and {\sc M.Z. Solomyak}, {\em Remarks on the spectral
shift function},  Zapiski Nauchn. Semin. LOMI {\bf27} (1972), 33--46 (Russian).

English transl.: J. Soviet Math. {\bf3} (1975), 408--419.

\bibitem[BS4]{BS4} {\sc M.S. Birman} and {\sc M.Z. Solomyak},
{\em Double Stieltjes operator integrals. III},
Problems of Math. Phys., Leningrad. Univ. {\bf6} (1973), 27--53 (Russian).

%\bibitem[BS5]{BS0} {\sc M.Sh. Birman and M.Z. Solomyak}, {\it Spectral theory of
%selfadjoint operators in Hilbert space},
%Mathematics and its Applications (Soviet Series),
%D. Reidel Publishing Co., Dordrecht, 1987.

\bibitem[BS5]{BS} {\sc M.S. Birman} and {\sc M.Z. Solomyak},
{\em Tensor product of a finite number of spectral measures is always a spectral measure}, Integral Equations Operator Theory {\bf24} (1996),  179--187.

%\bibitem[BS7]{BS7} {\sc M.S. Birman} and {\sc M.Z. Solomyak}, {\em Double operator integrals
%in Hilbert space},  Int. Equat. Oper. Theory  {\bf47}  (2003), 131--168.

\bibitem[DK]{DK} {\sc Yu.L. Daletskii} and {\sc S.G. Krein}, {\em Integration and differentiation of
functions of Hermitian operators and application to the theory of perturbations} (Russian), Trudy Sem.
Functsion. Anal., Voronezh. Gos. Univ. {\bf1} (1956), 81--105.

%\bibitem[Ca]{Ca} {\sc F. W.Carroll} {\em Mathematical Notes: Separately Continuous Functions
%are Baire Functions}, Amer. Math. Monthly {\bf78:2} (1971), 175.

%\bibitem[F1]{F1}  {\sc Yu.B. Farforovskaya}, {\em  The connection of the Kantorovich-Rubinshtein metric for spectral resolutions of selfadjoint operators with functions of operators},
%Vestnik Leningrad. Univ.  {\bf19}  (1968), 94--97. (Russian).

\bibitem[F]{F2}  {\sc Yu.B. Farforovskaya}, {\em An example of a Lipschitzian function of selfadjoint
operators that yields a nonnuclear increase under a nuclear perturbation}.  Zap. Nauchn. Sem.
Leningrad. Otdel. Mat. Inst. Steklov. (LOMI)  {\bf30}  (1972), 146--153 (Russian).


%\bibitem[5]{F2}  {\sc Yu.B. Farforovskaya},
%{\em An estimate of the norm of $\mid f(B)-f(A)\mid $ for selfadjoint operators $A$ and $B$},
%Zap. Nauchn. Sem. Leningrad. Otdel. Mat. Inst.
%Steklov. (LOMI)  {\bf56}  (1976), 143--162 (Russian).

%\bibitem[FN1]{FN1} {\sc Yu.B. Farforovskaya} and {\sc L.N. Nikolskaya},
%{\em An inequality for commutators of normal operators}, Acta Sci. Math. (Szeged) {\bf71} (2005), 751--765.
%
%\bibitem[FN2]{FN2} {\sc Yu.B. Farforovskaya} and {\sc L.N. Nikolskaya}, {\em Operator H\"olderness of
%H\"older functions} (Russian), to appear in
% Algebra i Analiz.

%\bibitem[GK]{GK} I.C. Gohberg and M.G. Krein, {\it Introduction to the theory
%of linear nonselfadjoint operators in Hilbert space,} Nauka, Moscow, 1965.

English transl.: Amer. Math. Soc., Providence, RI, 1969.

\bibitem[JW]{JW} {\sc B.E. Johnson} and {\sc J.P. Williams}, {\em The range of a normal derivation},
Pacific J. Math. {\bf58} (1975), 105--122.

\bibitem[Ka]{Ka}  {\sc T. Kato}, {\em Continuity of the map $S\mapsto \mid S\mid $ for linear operators},
Proc. Japan Acad.  {\bf49}  (1973), 157--160.

\bibitem[KS]{KS} {\sc E. Kissin} and {\sc V. S. Shulman}, {\em Classes of operator-smooth functions.
I Operator-lipschitz functions}, Proc. Edinburgh Math. Soc. {\bf 48} (2005), 151--173.

\bibitem[KPSS]{KPSS} {\sc E. Kissin, D. Potapov, V. S. Shulman} and {\sc F. Sukochev}, {\it Operator smoothness in Schatten norms for functions of several variables: Lipschitz conditions, differentiability and unbounded derivations},  Proc. Lond. Math. Soc. (3) {\bf105} (2012), 661--702.

\bibitem[Kr]{Kr} {\sc M.G. Krein}, {\em On a trace formula in perturbation theory},
Mat. Sbornik {\bf33} (1953), 597--626 (Russian).

%\bibitem[Kr2]{Kr2} {\sc M.G. Krein}, {\em Some new studies in the theory of perturbations of self-adjoint operators}, First Math. Summer School, Part I (Russian), 103-–187, Naukova Dumka, Kiev, 1964.

\bibitem[L]{L} {\sc I.M. Lifshitz}, {\em On a problem in perturbation theory
connected with quantum statistics}, Uspekhi Mat. Nauk {\bf7} (1952), 171--180 (Russian).

%\bibitem[L]{L}
%{\sc B.~Ya. Levin}, {\em Lectures on entire functions},
%Translation of Math. Monogr., vol. 150, 1996.

\bibitem[Mc]{Mc}
{\sc A. McIntosh}, {\em Counterexample to a question on commutators},
Proc. Amer. Math. Soc. {\bf29} (1971) 337--340.

\bibitem[Pee]{Pee} {\sc J. Peetre},
{\em New thoughts on Besov spaces}, Duke Univ. Press., Durham, NC, 1976.


\bibitem[Pe1]{Pe1} {\sc V.V.Peller}, {\em Hankel operators of class ${\bf S}_{p}$
and their applications (rational approximation, Gaussian processes,
the problem of majorizing operators)}, Mat. Sbornik,
{\bf 113} (1980), 538-581.

English Transl. in Math. USSR Sbornik, {\bf 41}
(1982), 443-479.


\bibitem[Pe2]{Pe2} {\sc V.V. Peller},
{\em Hankel operators in the theory of perturbations of unitary and self-adjoint operators},
Funktsional. Anal. i Prilozhen. {\bf19:2}  (1985),
37--51 (Russian).
English transl.: Funct. Anal. Appl. {\bf19} (1985) , 111--123.

%\bibitem[Pe]{Pe0} {\sc V.V. Peller}, {\em For which $f$ does $A-B\in{\bf S}_{p}$
%imply that $f(A)-f(B)\in{\bf S}_{p}$?}, Operator Theory, Birkh\"{a}user,
%{\bf 24} (1987), 289-294.


\bibitem[Pe3]{Pe3} {\sc V.V. Peller},
{\em Hankel operators in the perturbation theory of of unbounded self-adjoint operators}.
Analysis and partial differential equations,  529--544,
Lecture Notes in Pure and Appl. Math., {\bf122}, Dekker, New York, 1990.

\bibitem[Pe4]{Pe4} {\sc V.V. Peller}, {\em Hankel operators and their applications,}
Springer-Verlag, New York, 2003.

\bibitem[Pe5]{Pe5} {\sc V.V. Peller}, {\em Multiple operator integrals in perturbation theory}, Bull. Math. Sci.,

%\bibitem[Pe5] {Pe5}  {\sc V.V. Peller}, {\em Differentiability of functions of contractions},
%In: Linear and complex analysis, AMS Translations, Ser. 2 {\bf226}
%(2009), 109--131, AMS, Providence.

\bibitem[Pi]{Pi} {\sc G. Pisier}, {\em Similarity problems and completely bounded maps},
Second, expanded edition. Includes the solution to ``The Halmos problem''. Lecture Notes in Mathematics,
1618. Springer-Verlag, Berlin, 2001.

%\bibitem[Ru]{Ru} {\sc W. Rudin}. {\em Lebesgue's first theorem},
%Mathematical analysis and applications, Part B, pp. 741–-747, Adv. in Math. Suppl. Stud., 7b,
%Academic Press, New York-London, 1981.

%\bibitem[SKM]{SKM} {\sc St.G. Samko, A.A. Kilbas, O.I  Marichev},
%{\em Fractional integrals and derivatives: theory and applications},
%(Russian), Minsk, Nauka i Tekhnika, 1987.
%
%English Translation: Gordon and Breach, New York, NY, 1993.

%\bibitem[S]{S} {\sc E.M. Stein}, {\em Singular integrals and differentiability properties of functions},
%Princeton Math. Ser., No. 30, Princeton Univ. Press, Princeton, NJ, 1970,
%

%\bibitem[St]{S} {\sc V.V. Sten'kin}, {\em Multiple operator integrals},
%Izv. Vyssh. Uchebn. Zaved. Matematika {\bf4 (79)} (1977), 102--115 (Russian).

%English transl.: Soviet Math. (Iz. VUZ) {\bf21:4} (1977), 88--99.

%\bibitem[W]{W} {\sc H. Widom}, {\it When are differentiable functions differentiable?}, In: Linear and
%Complex Analysis Problem Book, Lect. Notes Math. {\bf 1043} (1984), 184--188.



%\bibitem[SNF]{SNF} {\sc B. Sz.-Nagy and C. Foias},
%{\it Harmonic analysis of operators on Hilbert
%space,} Akad\'{e}miai Kiad\'{o}, Budapest, 1970.


%\bibitem[SS]{SS} {\sc M.Z. Solomyak} and {\sc V.V. Sten'kin},
%{\em A certain class of multiple operator Stieltjes integrals},
%Problems of Math. Anal., No. 2: Linear Operators and Operator Equations (Russian), 122--134.
%Izdat. Leningrad. Univ., Leningrad, 1969.

%\bibitem[T]{T} {\sc H. Triebel}, {\em Theory of function spaces,}
%Monographs in Mathematics, {\bf78}, Birkh\"auser Verlag, Basel, 1983.
%
%\bibitem[W]{W} {\sc G.N. Watson}, {\em
%A treatise on the theory of Bessel functions,}
%Cambridge Mathematical Library. Cambridge: Cambridge Univ. Press. 1995.
%

\end{thebibliography}

\documentclass[11pt,reqno]{amsart}
\usepackage{amssymb,amscd,amsbsy}
\usepackage{amssymb,amscd,amsbsy,mathrsfs}
\newcommand{\lb}{\linebreak}

\newcommand{\s}{\sigma}

\newcommand{\f}{\varphi}

\newcommand{\D}{\Delta}
\renewcommand{\L}{\Lambda}

\newcommand{\E}{{\mathscr E}}

\newcommand{\h}{{\mathscr H}}

\newcommand{\X}{{\mathscr X}}
\newcommand{\Y}{{\mathscr Y}}

\newcommand{\R}{{\Bbb R}}

\newcommand{\bs}{\boldsymbol}

\newcommand{\m}{{\boldsymbol m}}
\newcommand{\bS}{{\boldsymbol S}}

\newcommand{\rf}[1]{(\ref{#1})}

\newcommand{\df}{\stackrel{\mathrm{def}}{=}}

\newcommand{\trace}{\operatorname{trace}}

\newcommand{\eeq}{\end{equation}}
\newcommand{\beq}{\begin{equation}}
\newcommand{\bay}{\begin{eqnarray}}
\newcommand{\ba}{\begin{align*}}
\newcommand{\ea}{\end{align*}}
\newcommand{\ey}{\end{eqnarray}}
\newcommand{\bey}{\begin{eqnarray*}}
\newcommand{\eey}{\end{eqnarray*}}

\newcommand{\be}{\infty}

\newcommand{\bl}{\blacksquare}

\newcommand{\Pf}{{\bf Proof. }}

\newtheorem{thm}{\hspace{\parindent}Theorem}[section]

\theoremstyle{remark}

\newtheorem*{rem*}{Remark}

\newcommand{\OL}{{\rm OL}}

\newcommand\Li{{\rm Lip}}
\newcommand\fM{\frak M}

\newcommand\dg{\frak D}

\begin{document}

\newcommand{\vse}{\vspace{.2in}}
\numberwithin{equation}{section}

\title{The Lifshits--Krein trace formula\\ and operator Lipschitz functions}
\author{V.V. Peller}
\thanks{the author is partially supported by NSF grant DMS 1300924}

\begin{abstract}
We describe the maximal class of functions $f$ on the real line, for which 
the Lifshitz--Krein trace formula $\trace(f(A)-f(B))=\int_\R f'(s)\bs{\xi}(s)\,ds$
holds for arbitrary self-adjoint operators $A$ and $B$ with $A-B$ in the trace class $\bS_1$. We prove that this class of functions coincide with the class of operator Lipschitz functions. 
\end{abstract}

\maketitle

\tableofcontents
\normalsize

\

%
%
%
%

\setcounter{section}{0}
\section{\bf Introduction}
\setcounter{equation}{0}
\label{In}

\

The purpose of this paper is to describe the class functions, for which the Lifshitz--Krein trace formula holds. The Lifshitz--Krein trace formula plays a significant role in perturbation theory. It was discovered by Lifshits \cite{L} in a special case and by Krein \cite{Kr} in the general case. This formula allows one to compute the trace of the difference
$f(A)-f(B)$ of a function $f$ of an unperturbed self-adjoint operator $A$ and a perturbed self-adjoint operator $B$
provided the perturbation $B-A$ belongs to trace class $\bS_1$.
M.G. Krein proved that for each such pair there exists a unique function $\bs{\xi}$ in $L^1(\R)$ such that for every function $f$ whose derivative is
the Fourier transform of and $L^1$ function, the operator $f(A)-f(B)$ belongs to $\bS_1$ and the following trace formula holds:
\bay
\label{LKf}
\trace\big(f(A)-f(B)\big)=\int_\R f'(s)\bs{\xi}(s)\,ds
\ey
(see \cite{Kr}). The function $\bs{\xi}$ is called the {\it spectral shift function associated with the pair} $(A,B)$. Clearly, the right-hand side of \rf{LKf} makes sense for arbitrary Lipschitz function $f$. In this connection Krein asked the question of whether it is true that for an arbitrary Lipschitz function $f$, the operator $f(A)-f(B)$ is in $\bS_1$ and trace formula \rf{LKf} holds.
{\it It turns out that this is false}. In \cite{F2} Farforovskaya gave an example of self-adjoint operators $A$ and $B$ with $A-B\in\bS_1$ and a Lipschitz function $f$ such that $f(A)-f(B)\not\in\bS_1$. 

Later it was shown in \cite{Pe2} and \cite{Pe3} that formula \rf{LKf} holds, whenever $A$ and $B$ are self-adjoint operators with $A-B\in\bS_1$ and $f$ belongs to the Besov space $B_{\be,1}^1(\R)$ (we refer the reader to \cite{Pee} and \cite{Pe4} for an introduction to Besov classes). Necessary conditions are also obtained in \cite{Pe2} and \cite{Pe3}. In particular, it was shown in \cite{Pe2} and \cite{Pe3} that 
if $f(A)-f(B)\in\bS_1$ whenever $A$ and $B$ are self-adjoint operators with $A-B\in\bS_1$, then $f$ locally belongs to the Besov space $B_{1,1}^1(\R)$. Note that those necessary conditions were deduced from the description of trace class 
Hankel operators \cite{Pe1} (see also \cite{Pe4}).

The main objective of this paper is to describe the class of functions $f$, for which
trace formula \rf{LKf} holds for arbitrary self-adjoint operators $A$ and $B$ with $A-B\in\bS_1$.

It is well known (see e.g. \cite{AP}) that for a function $f$ on $\R$, the following properties are equivalent:

(i) {\it there exists a positive number $C$ such that 
\bay
\label{OLn}
\|f(A)-f(B)\|\le C\|A-B\|
\ey
for all bounded self-adjoint operators $A$ and $B$;}

(ii) {\it there exists a positive number $C$ such that inequality {\em\rf{OLn}} holds,
whenever $A$ and $B$ are (not necessarily bounded) self-adjoint operators such that $A-B$ is bounded;}

(iii) {\it there exists a positive number $C$ such that 
\bay
\label{OLyad}
\|f(A)-f(B)\|_{\bS_1}\le C\|A-B\|_{\bS_1}
\ey
for all bounded self-adjoint operators $A$ and $B$ with $A-B\in\bS_1$;}

(iv) {\it there exists a positive number $C$ such that inequality {\em\rf{OLyad}} holds,,
whenever $A$ and $B$ are (not necessarily bounded) self-adjoint operators such that 
$A-B\in\bS_1$;}

(v) $f(A)-f(B)\in\bS_1${\it, whenever $A$ and $B$ are (not necessarily bounded) self-adjoint operators such that $A-B\in\bS_1$.}

Note that the minimal value of the constant $C$ is the same in (i)--(iv).

Functions satisfying (i) are called {\it operator Lipschitz}. We denote by ${\rm OL}(\R)$ the space of operator Lipschitz functions on $\R$. For 
$f\in{\rm OL}(\R)$, we define its quasi norm $\|f\|_{\rm OL}$ as the infimum of all constants $C$, for which inequality \rf{OLn} holds.
In other words,
\begin{align*}
\|f\|_{\rm OL}&=
\sup\left\{\frac{\|f(A)-f(B)\|_{\bS_1}}{\|A-B\|_{\bS_1}}:
~A~\mbox{ and }~B~\mbox{ are self-adjoint},~A-B~\mbox{ is bounded}\right\}\\[.2cm]
&=\sup\left\{\frac{\|f(A)-f(B)\|}{\|A-B\|}:
~A~\mbox{ and }~B~\mbox{ are self-adjoint},~A-B\in\bS_1\right\}.
\end{align*}

It was shown in \cite{JW} that operator Lipschitz functions are differentiable everywhere on $\R$. Note that this implies that the function $x\mapsto|x|$
is not operator Lipschitz, the fact established earlier in \cite{Mc} and \cite{Ka}. On the other hand, an operator Lipschitz function does not have to be continuously differentiable; in particular, the function $x\mapsto x^2\sin x^{-1}$
is operator Lipschitz, see \cite{KS}.

For a differentiable function $f$ on $\R$, we consider the divided difference
$\dg f$ defined by
\bay
\label{razdraz}
(\dg f)(x,y)=\left\{\begin{array}{ll}\frac{f(x)-f(y)}{x-y},&x\ne y\\[.2cm]
f'(x),&x=y.\end{array}
\right.
\ey
It turns out (see e.g., \cite{AP}) that a differentiable function on $\R$ is operator Lipschitz if and only if the divided difference $\dg f$ is a Schur multiplier (see \S\,\ref{dois}) for the definition.

The main purpose of this paper is to prove that {\it the condition $f\in{\rm OL}(\R)$
is not only a necessary condition for the Lifshits--Krein trace formula {\em\rf{LKf}}
to hold for arbitrary self-adjoint operators $A$ and $B$ with $A-B\in\bS_1$, but is also sufficient.} This will be proved in \S\,\ref{self}.

In \S~\ref{dois} We define double operator integrals and Schur multipliers.
In \S~\ref{HSdif} we state a result of \cite{KPSS} on the differentiability of the function $t\mapsto f(A+tK)-f(A)$ in the Hilbert--Schmidt norm. We state a characterization of the space of Schur multipliers in terms of Haagerup tensor products in \S~\ref{Haag}.
Finally, in \S~\ref{sleddoi} we obtain a formula for the trace of double operator integrals.


%
%
%

\

\section{\bf Double operator integrals and Schur multipliers}
\setcounter{equation}{0}
\label{dois}

\

Double operator integrals appeared in the paper \cite{DK} by Daletskii and S.G. Krein. 
Later the beautiful theory of double operator integrals was created by Birman and Solomyak in \cite{BS1}, \cite{BS2}, and \cite{BS4}. 

Let $(\X,E_1)$ and $(\Y,E_2)$ be spaces with spectral measures $E_1$ and $E_2$
on a Hilbert space $\h$ and let $\Phi$ be a bounded measurable function on $\X\times\Y$. Double operator integrals are expressions of the form
\bay
\label{doi}
\int\limits_\X\int\limits_\Y\Phi(x,y)\,d E_1(x)T\,dE_2(y).
\ey
Birman and Solomyak starting point is the case when $T$ belongs to the Hilbert--Schmidt class $\bS_2$. For a bounded measurable function on $\Phi$ on $\X\times\Y$ and an operator $T$ of class $\bS_2$, consider the spectral measure $\E$ whose values are orthogonal
projections on the Hilbert space $\bS_2$, which is defined by
$$
\E(\L\times\D)T=E_1(\L)TE_2(\D),\quad T\in\bS_2,
$$
$\L$ and $\D$ being measurable subsets of $\X$ and $\Y$. 
It was shown in \cite{BS} that $\E$ extends to a spectral measure on
$\X\times\Y$. For a bounded measurable function $\Phi$ on $\X\times\Y$, the double operator integral \rf{doi} is defined by
$$
\int\limits_\X\int\limits_\Y\Phi(x,y)\,d E_1(x)T\,dE_2(y)\df
\left(\,\,\int\limits_{\X\times\Y}\Phi\,d\E\right)T.
$$
Clearly,
$$
\left\|\int\limits_\X\int\limits_\Y\Phi(x,y)\,dE_1(x)T\,dE_2(y)\right\|_{\bS_2}
\le\|\Phi\|_{L^\be}\|T\|_{\bS_2}.
$$
If
$$
\int\limits_\X\int\limits_\Y\Phi(x,y)\,d E_1(x)T\,dE_2(y)\in\bS_1
$$
for every $T\in\bS_1$, we say that $\Phi$ is a {\it Schur multiplier of $\bS_1$ associated with
the spectral measures $E_1$ and $E_2$}. {\it We denote by $\fM(E_1,E_2)$ the space of
Schur multipliers of $\bS_1$ with respect to $E_1$ and $E_2$.} The norm 
$\|\Phi\|_{\fM(E_1,E_2)}$ of $\Phi$ in the space $\fM(E_1,E_2)$ is, by definition, the norm of the linear transformer
$$
T\mapsto\int\limits_\X\int\limits_\Y\Phi(x,y)\,d E_1(x)T\,dE_2(y)
$$
on the class $\bS_1$.

If $\Phi\in\fM(E_1,E_2)$, one can define by duality double operator integrals of the form \rf{doi} for an arbitrary bounded linear operator $T$. However, we do not need this in this paper.

We are going to discuss briefly in \S~\ref{Haag} characterizations of Schur multipliers.

Birman and Solomyak proved in \cite{BS4} that if $f$ is a Lipschitz function and $A$ and $B$ are not necessarily bounded self-adjoint operators with $A-B\in\bS_2$, then 
\bay
\label{razS2}
f(A)-f(B)=\iint\limits_{\R\times\R}\frac{f(x)-f(y)}{x-y}\,dE_A(x)(A-B)\,dE_B(y).
\ey
Note that for an arbitrary Lischitz function $f$, the divided difference $\dg f$
is not always naturally defined on the diagonal. However, we can define $\dg f$
on the diagonal by an arbitrary bounded measurable function and the right-hand side
of \rf{razS2} does not depend on the values on the diagonal. It follows from \rf{razS2} that
$$
\|f(A)-f(B)\|_{\bS_2}\le\|f\|_{\Li}\|A-B\|_{\bS_2},
$$
where the Lipschitz (semi)norm $\|f\|_{\Li}$ of $f$ is, by definition,
$$
\sup\left\{\frac{|f(x)-f(y)|}{|x-y|}:~x,~y\in\R,~x\ne y\right\}.
$$

On the other hand, if $A$ and $B$ are not necessarily bounded self-adjoint operators with $A-B\in\bS_1$ and $f$ is an operator Lipschitz function, then
\bay
\label{razS1}
f(A)-f(B)=\iint\limits_{\R\times\R}\big(\dg f\big)(x,y)\,dE_A(x)(A-B)\,dE_B(y).
\ey
(see \cite{BS4}). Here the divided difference $\dg f$ is defined by \rf{razdraz}.
It follows from \rf{razS1} that
$$
\|f(A)-f(B)\|_{\bS_1}\le\|f\|_{\OL}\|A-B\|_{\bS_1}.
$$

\

\section{\bf Differentiability in the Hilbert--Schmidt norm}
\setcounter{equation}{0}
\label{HSdif}

\

Suppose that $A$ and $B$ are not necessarily bounded self-adjoint operators on Hilbert space such that $A-B\in\bS_2$. Consider the parametric family $A_t$, $0\le t\le1$, defined by $A_t=A+tK$, where $K\df B-A$. We need the following result of \cite{KPSS}, Theorem 7.18:

{\it Suppose that $f$ is a Lipschitz function on $\R$ that is differentiable at every point of $\R$. Then the function $s\mapsto f(A_s)-f(A)$ is differentiable on $[0,1]$ in the Hilbert--Schmidt norm and
$$
\frac{d}{ds}\big(f(A_s)-f(A)\big)\Big|_{s=t}
=\iint\limits_{\R\times\R}(\dg f)(x,y)\,dE_t(x)K\,dE_t(y).
$$
}

\

\section{\bf Schur multipliers and Haagerup tensor products}
\setcounter{equation}{0}
\label{Haag}

\

Let $(\X,E_1)$ and $(\Y,E_2)$ be spaces with spectral measures $E_1$ and $E_2$
on Hilbert space. There are several characterizations of the class $\fM(E_1,E_2)$ 
of Schur multipliers, see \cite{Pe2}, \cite{Pi}, \cite{AP}. We need the following characterization in terms of the Haagerup tensor product of $L^\be$ spaces:

{\it Let $\Phi$ be a measurable function on $\X\times\Y$. Then $\Phi\in\fM(E_1,E_2)$ if and only if $\Phi$ belongs to the Haagerup tensor product 
$L^\be(E_1)\!\otimes_{\rm h}\!L^\be(E_2)$, i.e., $\Phi$ admits a representaion
$$
\Phi(x,y)=\sum_n\f_n(x)\psi_n(y),
$$
where $\f_n\in L^\be(E_1)$, $\psi_n\in L^\be(E_2)$, and}
$$
\sum_n|\f_n|^2\in L^\be(E_1)\quad\mbox{and}\quad\sum_n|\psi_n|^2\in L^\be(E_2).
$$

Suppose now that $E_1$ and $E_2$ are Borel spectral measures on locally compact topological spaces $X$ and $Y$. In this case the following result holds (see \cite{AP}):

{\it Let $\Phi$ be a function on $\X\times\Y$ that is continuous in each variable.
Then \lb$\Phi\in\fM(E_1,E_2)$ if and only if it belongs to the Haagerup tensor product
$C_{\rm b}(\X)\!\otimes_{\rm h}\!C_{\rm b}(\Y)$ of the spaces of bounded continuous functions on $\X$ and $\Y$, i.e., $\Phi$ admits a representation 
$$
\Phi(x,y)=\sum_n\f_n(x)\psi_n(y),
$$
where $\f_n\in C_{\rm b}(\X)$, $\psi_n\in C_{\rm b}(\Y)$ and the functions
$$
\sum_n|\f_n|^2\quad\mbox{and}\quad\sum_n|\psi_n|^2
$$
are bounded.}

\

\section{\bf The trace of double operator integrals}
\setcounter{equation}{0}
\label{sleddoi}

\

Let $T$ be a trace class operator on Hilbert space and let
$E$ be a spectral measure on a $\s$-algebra of subsets of a set $\X$.
If $\Phi$ is a Schur multiplier, i.e., $\Phi\in\fM(E,E)$, then the double operator integral
$$
\iint\Phi(x,y)\,dE(x)T\,dE(y)
$$
belongs to $\bS_1$. Let us compute its trace. In \cite{BS4} the following trace formula
was found:
\bay
\label{BStf}
\trace\left(\iint\Phi(x,y)\,dE(x)T\,dE(y)\right)=\int\Phi(x,x)\,d\mu(x),
\ey
where $\mu$ is the complex measure defined by
$$
\mu(\D)=\trace\big(TE(\D)\big).
$$

The problem is how we can interpret the function $x\mapsto\Phi(x,x)$
for functions $\Phi$ in $\fM(E,E)$. In 
\cite{Pe5} the following justification of formula \rf{BStf} was given. We can define the trace ${\mathscr T}\Phi$ of a function $\Phi$ in $\fM(E,E)$  on the diagonal by the formula
$$
({\mathscr T}\Phi)(x)\df\sum_n\f_n(x)\psi_n(x),
$$
where 
\bay
\label{Htpre}
\Phi(x,y)=\sum_n\f_n(x)\psi_n(y)
\ey
is a representation of $\Phi$ as an element of the Haagerup tensor product 
$L^\be(E)\otimes_{\rm h}L^\be(E)$, i.e.,
\bay
\label{Htner}
\sum_m|\f_n|^2\in L^\be(E)\quad\mbox{and}\quad\sum_m|\psi_n|^2\in L^\be(E).
\ey
Clearly, the trace of $\Phi\in\fM(E,E)$ on the diagonal belongs to $L^\be(E)$.
Then formula \rf{BStf} holds if $\Phi(x,x)$ is understood as $({\mathscr T}\Phi)(x)$,
see \cite{Pe5}, \S~1.1.

Suppose now that $E$ is a Borel spectral measure on a locally compact topological space 
$\X$ and $\Phi$ is a function on $\X\times\X$ that is continuous in each variable.
As we have mentioned in \S~\ref{dois}, $\Phi$ admits a representation of the form
\rf{Htpre}, in which the functions $\f_n$ and $\psi_n$ satisfy \rf{Htner} and are continuous functions on $\X$. It is easy to see that in this case 
$({\mathscr T}\Phi)(x)=\Phi(x,x)$, $x\in\X$. In other words, the following theorem holds:

\begin{thm}
\label{sleddvoi}
Let $E$ be a spectral measure on a locally compact topological space $\X$. Suppose that 
$\Phi$ is a function of class $\fM(E,E)$. If $\Phi$ is continuous in each variable, then
formula {\em\rf{BStf}} holds for an arbitrary trace class operator $T$.
\end{thm}

\

\section{\bf The Lifshits--Krein trace formula for \\ arbitrary operator Lipschitz functions}
\setcounter{equation}{0}
\label{self}

\

Suppose that $A$ and $B$ are self-adjoint operators on Hilbert space such that
$A-B\in\bS_1$. Let $\bs\xi$ be the spectral shift function associated with the pair
$(A,B)$. As we have mentioned in \S~\ref{dois}, for an arbitrary operator Lipschitz function $f$
on $\R$, the operator $f(A)-f(B)$ belongs to trace class. The following theorem is the main result of the paper.

\begin{thm}
\label{glavn}
Let $f\in\OL(\R)$. Then
$$
\trace\big(f(A)-f(B)\big)=\int_\R f'(s)\bs{\xi}(s)\,ds.
$$
\end{thm}

To prove the theorem, we are going to use an approach of Birman and Solomyak in \cite{BSsl} to the Lifshits--Krein trace formula. In \cite{BSsl} they used their approach under more restrictive assumptions on $f$.

\medskip

\Pf Obviously, $f$ is a Lipschitz function. As we have mentioned in the introduction, $f$ is a differentiable function at every point of $\R$ (but not necessarily continuously differentiable!). 
Put $K\df B-A$.
Consider the parametric family $\{A_t\}_{0\le t\le1}$, $A_t\df A+tK$. Then $A_0=A$ and $A_1=B$. The operator $K$ obviously belongs to the Hilbert--Schmidt class $\bS_2$. As we have mentioned in \S~\ref{HSdif}, the function
\mbox{$t\mapsto f(A_t)-f(A)$} is differentiable in the Hilbert--Schmidt norm and
$$
Q_t\df
\frac{d}{ds}\big(f(A_s)-f(A)\big)\Big|_{s=t}=
\iint\limits_{\R\times\R}\frac{f(x)-f(y)}{x-y}\,dE_t(x)K\,dE_t(y)\in\bS_2,
$$
where $E_t$ is the spectral measure of $A_t$.

On the other hand, since the divided difference $\dg f$ is a Schur multiplier of of $\bS_1$ (see the introduction), it follows that
$$
Q_t\in\bS_1,\quad 0\le t\le1,\quad\mbox{and}\quad\sup_{t\in[0,1]}\|Q_t\|_{\bS_1}<\be.
$$

We have
$$
f(A)-f(B)=-\int_0^1Q_t\,dt,
$$
where the integral on the right is understood in the sense of Bochner in the space $\bS_1$. It follows that
$$
\trace\big(f(A)-f(B)\big)=-\int_0^1\trace(Q_t)\,dt.
$$

Since the function $f$ is differentiable everywhere, the divided difference $\dg f$ 
is continuous in each variable. By Theorem \ref{sleddvoi},
$$
\trace Q_t=\int_\R f'(x)\,d\nu_t(x),
$$
where the signed measure $\nu_t$ is defined by
$$
\nu_t(\D)\df \trace\big(E_t(\D)K)\quad\mbox{for a Borel subset}\quad\D\quad\mbox{of}\quad\R.
$$ 

We identify here the space of complex Borel measures on $\R$ with the dual space to the Banach space of continuous functions on $\R$ with zero limit at infinity. Then the function $t\mapsto \nu_t$ is continuous in the weak-$*$ topology on the space of complex Borel measures. Indeed, if $h$ is continuous on $\R$ and $\lim_{|x|\to\be}h(x)=0$, then
$$
\int h\,d\nu_t=\trace\big(h(A_t)K\big).
$$ 
The function $t\mapsto h(A_t)$ is a continuous function on $[0,1]$ in the operator norm; this follows from the fact that $h$ is an operator continuous function (see \cite{AP2}, \S~8). Thus 
the function $t\mapsto\trace\big(h(A_t)K\big)$ is continuous.

Therefore we can define the signed Borel measure $\nu$ on $\R$ by
$$
\nu=\int_0^1\nu_t\,dt.
$$
It follows that
$$
\trace\big(f(A)-f(B)\big)=-\int_\R f'\,d\nu.
$$

On the other hand, for smooth functions $g$ with compact support,
$$
\trace\big(g(A)-g(B)\big)=\int_\R g'\bs{\xi}\,d\m,
$$
where $\bs{\xi}$ is the spectral shift function. 

It follows that $\nu$ is absolutely continuous with respect to Lebesgue measure and
\lb$d\nu=-\bs{\xi}d\m$. $\bl$

\medskip

Theorem \ref{glavn} implies the following result:

\begin{thm}
Let $f$ be an operator Lipschitz function and let $A$ and $B$ be self-adjoint operators such that $A-B\in\bS_1$. Then the function
$$
t\mapsto\trace\big(f(A-tI)-f(B-tI)\big),\quad t\in\R,
$$
is continuous on $\R$.
\end{thm}

\Pf Consider the function $f_t$, $t\in\R$, defined by $f_t(x)\df f(x-t)$. Let $\bs{\xi}$
be the spectral shift function associated with $(A,B)$. It is easy to see that
$$
\trace\big(f(A+tI)-f(B+tI)\big)=\trace\big(f_t(A)-f_t(B)\big)
=\int_\R f'(x-t)\bs{\xi}(x)\,d\m(x),
$$
which depends on $t$ continuously. $\bl$

\

\

\

\footnotesize
\noindent
V.V. Peller \\
Department of Mathematics  \\
Michigan State University\\
East Lansing Michigan 48824\\

\end{document}